OpenStudy (anonymous):

calculate the mass units of .25 mol of carbon -12 atoms

OpenStudy (anonymous):

???

OpenStudy (xishem):

\[\frac{0.25mol(^{12}C)}{1}*\frac{6.022*10^{23}atoms(^{12}C)}{1mol(^{12}C)}*\frac{12u}{1atom(^{12}C)}=1.8*10^{24}u\]Sorry. I had posted the solution and then deleted it, because there was a mistake in my old post.

OpenStudy (anonymous):

no worries <3

OpenStudy (anonymous):

mass=36gram